# Understanding Reader Perception Shifts upon Disclosure of AI Authorship

### v1

@@ Abstract.
As AI writing support becomes ubiquitous, how disclosing its use affects reader perception remains a critical, underexplored question. We conducted a study with 261 participants to examine how revealing varying levels of AI involvement shifts author impressions across six distinct communicative acts. Our analysis of 990 responses shows that disclosure generally erodes perceptions of trustworthiness, caring, competence, and likability, with the sharpest declines in social and interpersonal writing. A thematic analysis of participants’ feedback links these negative shifts to a perceived loss of human sincerity, diminished author effort, and the contextual inappropriateness of AI. Conversely, we find that higher AI literacy mitigates these negative perceptions, leading to greater tolerance or even appreciation for AI use. Our results highlight the nuanced social dynamics of AI-mediated authorship and inform design implications for creating transparent, context-sensitive writing systems that better preserve trust and authenticity.
@@ 1. Introduction
Artificial intelligence (AI) systems have reached a level of writing fluency comparable to human authors leading to their widespread adoption across professional, creative, and interpersonal contexts. While such tools can improve clarity and expression, they also introduce new social challenges. Readers often perceive AI-generated texts as less empathetic or authentic (Rubin et al., 2025), and disclosing AI authorship can degrade perceptions of authenticity and trust (Jakesch et al., 2019; Liu et al., 2022; Schilke and Reimann, 2025). This tension arises because writing is more than information transfer. It can also be a social act that conveys an author’s intention, effort, and emotion to readers. When an AI actively contributes to the writing process, this fundamental social meaning of authorship becomes ambiguous, complicating how a text and its author are perceived.
@@
While previous work in AI-mediated communication has outlined its ethical and relational challenges (Hancock et al., 2020), empirical studies have so far examined reader perceptions with limited granularity or control over AI authorship. Much of the existing research compares purely manual writing to AI-assisted text in specific, narrow contexts, often treating writing as a monolithic activity (Jain et al., 2024; Liu et al., 2022; Schilke and Reimann, 2025). This overlooks the nuanced nature of writing, which serves a wide spectrum of communicative goals. As articulated in frameworks like Berge et al.’s Wheel of Writing, writing can be used to persuade, to interact socially, to reflect internally, or to imagine creatively (Berge et al., 2016). The social expectations for an apology letter, for example, are vastly different from those for a technical manual. As AI writing support becomes deeply embedded in these varied contexts, the boundary between helpful assistance and core authorship blurs. Furthermore, readers’ attitudes toward writing co-authored by AI may vary depending on their familiarity with the technology. Previous studies have indicated that individuals who are more knowledgeable about AI tend to view its use as a pragmatic choice rather than a lack of competence (Mahmud et al., 2024). This raises critical, unanswered questions: How do readers’ perceptions shift based on the degree of AI authorship across these different communication purposes? Furthermore, how do individual factors, such as readers’ AI literacy, moderate these perceptions? Addressing these questions is essential for designing AI authoring tools that are not only effective but also socially acceptable and trustworthy.
@@
To address this research gap, we conducted a user study with 261 participants to examine how readers’ perceptions of a fictitious author change upon the disclosure of AI authorship across 6 acts of writing (Berge et al., 2016). We found consistent negative shifts in perceived trustworthiness, caring, competence, and likability, especially in emotional and interpersonal writing. We also found that participants with higher AI literacy exhibited greater tolerance toward and even appreciation of AI authorship. Our results reveal the complex social dynamics of AI authorship and inform the design of transparent, context-sensitive writing tools that preserve authenticity and trust. Our work positions AI writing assistance not as a purely technical problem of automation but as a relational design challenge. We argue that future interfaces must balance AI capabilities with the need to communicate human authorship, social context, and the perceived effort invested by the author.
@@
  • Design implications for future AI-mediated writing Based on our findings, we derive four design implications for future systems: 1) context-sensitive transparency that explains the purpose of AI authorship; 2) interfaces that preserve and communicate human effort and agency; 3) writing support adaptive to emotional and social contexts; and 4) reflective interfaces that foster AI literacy and calibrated trust.
@@ 2.1. AI-assisted Writing for Diverse Communicative Purposes
Prior research has explored how these tools support a range of writing goals. Mirowski et al. demonstrated that AI can be integrated into writers’ workflows to enhance creativity in screen- and playwriting (Mirowski et al., 2023). AI has also been leveraged in contexts where writing serves an interpersonal or social function. For instance, systems that suggest, generate, or refine email responses can help users convey positivity (Mieczkowski et al., 2021), politeness, and professionalism in workplace interactions (Miura et al., 2025), while AI-assisted dating profiles can enhance self-presentation (Barkallah et al., 2025). These applications demonstrate that AI’s role extends far beyond grammatical correction to shaping a writer’s tone, style, and communicative intent, aiding in both persuasion and social engagement (Hancock et al., 2020).
This expanding role, however, introduces social friction, particularly when AI’s contribution blurs the boundaries of authorship. Readers often question the authenticity and credibility of AI-generated content. For instance, Li et al. found that revealing AI’s role in writing essays lowered readers’ quality ratings (Li et al., 2024). Similarly, research on AI-mediated communication showed that online profiles perceived as AI-generated are less trusted (Jakesch et al., 2019). These concerns are particularly salient when writing serves a social purpose, such as persuasion or interaction, where impressions of the author are shaped by perceived authenticity. Recent studies on disclosure dynamics confirm this tension: while AI assistance can improve a text’s fluency, revealing its use can lead readers to judge the author as less sincere or capable (Hwang et al., 2025; Draxler et al., 2024).
AI-assisted writing thus presents a fundamental trade-off. On one hand, it empowers individuals, especially those with language barriers or weaker writing skills, to communicate more effectively (Ningrum et al., 2023; Noy and Zhang, 2023). On the other, the social meaning of authorship complicates its adoption. When readers know or suspect AI involvement, they may perceive the writing—and by extension, the author—as less authentic or trustworthy (Jakesch et al., 2019). Understanding how these perceptions vary across different communicative purposes is therefore critical for designing AI tools that successfully balance expressive power with the need to maintain credibility.
@@ 2.2. Reader Perceptions of AI Disclosure
Research on AI disclosure highlights its paradoxical effect on interpersonal trust. Several studies show that when readers learn AI contributed to a text, their trust in the human author declines even if the text’s quality is unchanged (Jakesch et al., 2019; Lermann Henestrosa and Kimmerle, 2024; Liu et al., 2022; Proksch et al., 2024; Schilke and Reimann, 2025). This ”transparency penalty” challenges the conventional assumption that openness builds trust (Sah et al., 2018). The magnitude of this trust erosion is not uniform and appears to be moderated by individual differences. For instance, Schilke and Reimann (2025) found that people with more positive attitudes toward AI show a smaller decline in trust upon disclosure. Furthermore, a user’s AI literacy plays a significant role in shaping their perceptions (Mahmud et al., 2024; Schiavo et al., 2024). AI-literate individuals tend to view its use as pragmatic, whereas less knowledgeable readers may see it as incompetence or laziness (Mahmud et al., 2024; Schiavo et al., 2024).
Complicating this landscape is the fact that people are increasingly unable to distinguish AI-generated text from human writing without explicit disclosure (Májovskỳ et al., 2024; Weber-Wulff et al., 2023), a finding that holds true even for those knowledgeable about LLMs (Dugan et al., 2023; Hakam et al., 2024; Porter and Machery, 2024). This creates a risk of mis-attribution, where readers may credit AI-generated text to a human, only to have their perception negatively altered if the AI’s role is later revealed. Despite the importance of this phenomenon, little research has systematically examined how varying degrees of AI involvement influence reader perception across different writing contexts.
Beyond trust, research also reveals mixed attitudes toward the social qualities of AI-generated text. Some studies find that AI is perceived as less caring, warm, or empathetic than humans, particularly in sensitive contexts like health communication (Rubin et al., 2025; Wang et al., 2025a). Conversely, other studies show that AI-generated content can be rated as equal or even superior to human-authored text, especially when qualities like efficiency and accuracy are valued (Ayers et al., 2023; Ovsyannikova et al., 2025).
One factor that may explain these divergent findings is people’s prior experience and familiarity with AI. Regular users may view AI authorship as beneficial (Mahmud et al., 2024), while less experienced individuals might perceive it as cold, inauthentic or inadequate (Li et al., 2024). Together, these findings suggest that AI disclosure has multifaceted effects, influencing not only trust in the author but also perceptions of warmth, competence, and empathy.
@@ 3. Method
To investigate how the disclosure of AI assistance may influence reader perception, we conducted a repeated-measures online study. Participants first rated their impressions of a text and then rated it again after we disclosed that a certain percentage of the text was generated or edited by AI. In contrast to prior work, our primary goal was to examine how perceptions shifted across systematically varying levels of disclosed AI contribution. To achieve this, we used a deception-based design: while all texts were entirely AI-generated (with minor human proofreading), we informed participants that only specific portions of each text were created or altered by AI. This approach allowed us to precisely control the perceived degree of AI involvement.
Our study was designed to answer the following research questions:
@@ 3.1. Experiment Design
We grounded our experimental design in the six fundamental acts of writing defined by Berge et al. (Berge et al., 2016): Convince, Interact, Reflect, Describe, Explore, and Imagine. As detailed in Table 1, these acts serve distinct communicative purposes. Convince, Interact, and Reflect are primarily person-oriented, directed at others, social relationships, and the self, respectively. On the other hand, Describe, Explore, and Imagine are object-oriented, focusing on organizing, developing, and creating knowledge.
@@
To imagine, narrate, create, and theorize Creation of textual words
To create our experimental stimuli, we generated three scenarios for each act, yielding 18 unique texts (Table 2). As the study was conducted with Japanese participants, all texts were in Japanese. The texts were first generated in English using GPT-4o  [^1] https://chatgpt.com and then translated into Japanese. A native Japanese-speaking author reviewed all translated texts and made minor corrections to ensure linguistic quality and naturalness. The final texts averaged 20 sentences and 662 Japanese characters in length. The full English and Japanese scripts are provided as supplementary material.
@@
  a tutorial or operating new equipment
@@
To manipulate the perceived level of AI involvement, we simulated various levels of AI authorship. For each text, a random proportion of sentences was labeled as having been generated or edited by AI (in reality, all texts were entirely AI-authored). The disclosed proportion varied from 0% to 100% in 10% increments across conditions, corresponding to approximately two sentences per increment (Figure 1). This design allowed us to model situations where a reader discovers that some portions of a text, from minor local edits to the entire document, were created with AI assistance.
@@ 3.2.1. Pre-disclosure Evaluation
After providing informed consent, participants were each assigned four of the 18 texts, representing different acts of writing, and were asked to read them carefully. For each text, they answered a questionnaire assessing their initial impressions of the (fictitious) author (Figure 1(a)). The questionnaire assessed five perception dimensions using 7-point Likert scales adapted from established instruments (27 statements in total). Trustworthiness (6 statements), Caring/Goodwill (6 statements), and Competence (6 statements) were adapted from McCroskey and Teven (McCroskey and Teven, 1999), and responses ranged from -3 (very negative) to +3 (very positive). Likability (7 statements) and the participant’s Desire for Future Interaction (2 statements), were adapted from Reysen (Reysen, 2005) and Sprecher (Sprecher, 2021). These were rated from -3 (Strongly disagree) to 3 (Strongly agree). All statements are included in Table 5 and 6 in the Appendix.
The same set of questions was used in both the pre- and post-disclosure phases. To ensure data quality, we included reverse-coded items and attention checks in both questionnaires.
@@ 3.2.2. Simulated AI Authorship Disclosure and Post-disclosure Evaluation
Following this disclosure, participants answered the same five questions to capture their revised impressions. They were also asked to provide a free-form text response explaining how and why their ratings had changed (if that was the case).
@@ 3.2.3. Post-experimental Questionnaire
After completing all four evaluations, participants filled out a final AI-literacy questionnaire to examine whether their familiarity and skills with AI tools influenced their perceptions. To measure this competency, we adapted three relevant sub-constructs from the Meta AI Literacy Scale (MAILS) (Carolus et al., 2023): “Apply AI”, “Detect AI”, and “AI Persuasion Literacy”. “Apply AI” measures an individual’s ability to use AI concepts for problem-solving, reflecting their proficiency and comfort with AI tools (Ng et al., 2022). “Detect AI” assesses the ability to distinguish technologies that use AI from those that do not, a core component of understanding what AI is (Long and Magerko, 2020; Wang et al., 2023). Finally, “AI Persuasion Literacy” refers to the ability to recognize and resist AI-driven influence on one’s decisions. Table 6 shows the items of each aspect of AI literacy.
@@
Figure 1. The user interfaces of the experiment: (a) pre-rating and (b) post-rating phases. The original interface was in Japanese and the text shown here is translated for presentation, but the layout and visual design are preserved.:

### v2

@@ Abstract.
As AI writing support becomes ubiquitous, the question of how disclosing its use affects reader perception remains critical and underexplored. We conducted a controlled study with 261 participants to examine how disclosing varying levels of AI involvement shifts perceptions of the author across six distinct communicative acts. Our analysis of 990 evaluations reveals that disclosure generally erodes perceived trustworthiness, caring, competence, and likability, with the most precipitous declines observed in social and interpersonal writing. A thematic analysis of participant feedback attributes these negative shifts to a perceived loss of human sincerity, diminished authorial effort, and the contextual inappropriateness of AI. Notably, however, we find that higher AI literacy mitigates these negative perceptions, leading to greater tolerance or even appreciation for AI assistance. Our results highlight the nuanced social dynamics of AI-mediated authorship and inform design implications for transparent, context-sensitive writing systems that better preserve trust and authenticity.
@@
Accepted at 31st International Conference on Intelligent User Interfaces (IUI ’26).
@@ 1. Introduction
Artificial intelligence (AI) systems have reached a level of writing fluency comparable to human authors leading to their widespread adoption across professional, creative, and interpersonal contexts. While these tools can enhance clarity and expression, they also introduce significant social frictions. Readers often perceive AI-generated texts as less empathetic or authentic (Rubin et al., 2025), and disclosing AI authorship can degrade perceptions of authenticity and trust (Jakesch et al., 2019; Liu et al., 2022; Schilke and Reimann, 2025). This tension arises because writing is more than mere information transfer. It is a social act that conveys an author’s intention, effort, and emotion to readers. When an AI actively contributes to the writing process, this fundamental social meaning of authorship becomes ambiguous, complicating how a text and its author are perceived.
@@
While previous work in AI-mediated communication has outlined these ethical and relational challenges (Hancock et al., 2020), empirical studies have so far examined reader perceptions with limited granularity or control over AI authorship. Much of the existing research compares purely manual writing to AI-assisted text in specific, narrow contexts, often treating writing as a monolithic activity (Jain et al., 2024; Liu et al., 2022; Schilke and Reimann, 2025). This overlooks the nuanced nature of writing, which serves a wide spectrum of communicative goals. As articulated in frameworks like Berge et al.’s Wheel of Writing, writing can be used to persuade, to interact socially, to reflect internally, or to imagine creatively (Berge et al., 2016). The social expectations for an apology letter, for example, are vastly different from those for a technical manual. As AI writing support becomes deeply embedded in these varied contexts, the boundary between helpful assistance and core authorship blurs. Furthermore, readers’ attitudes toward writing co-authored by AI may vary depending on their familiarity with the technology. Previous studies have indicated that individuals who are more knowledgeable about AI tend to view its use as a pragmatic choice rather than a lack of competence (Mahmud et al., 2024). This raises critical, unanswered questions: How do readers’ perceptions shift based on the degree of AI authorship across these different communication purposes? Furthermore, how do individual factors, such as readers’ AI literacy, moderate these perceptions? Addressing these questions is essential for designing AI authoring tools that are not only effective but also socially acceptable and trustworthy.
@@
To address this research gap, we conducted a controlled user study with 261 participants that examines how readers’ perceptions of a fictitious author change upon the disclosure of AI authorship across six distinct acts of writing (Berge et al., 2016). Our analysis revealed consistent negative shifts in perceived trustworthiness, caring, competence, and likability, especially in emotional and interpersonal writing. However, we also found that participants with higher AI literacy exhibited greater tolerance toward, and even appreciation of, AI authorship. Our results highlight the complex social dynamics of AI authorship and inform the design of transparent, context-sensitive writing tools that preserve authenticity and trust. We position AI writing assistance not merely as a technical problem of automation, but as a relational design challenge. We argue that future interfaces must balance AI capabilities with the need to communicate human authorship, social context, and the perceived effort invested by the author.
@@
  • Design implications for future AI-mediated writing: Based on our findings, we derive four design implications for future systems: 1) context-sensitive transparency that explains the purpose of AI authorship; 2) interfaces that preserve and communicate human effort and agency; 3) writing support adaptive to emotional and social contexts; and 4) reflective interfaces that foster AI literacy and calibrated trust.
@@ 2.1. AI-assisted Writing for Diverse Communicative Purposes
Prior research has explored how AI tools support a range of writing goals. Mirowski et al. demonstrated that AI can be integrated into writers’ workflows to enhance creativity in screen- and playwriting (Mirowski et al., 2023). AI has also been leveraged in contexts where writing serves an interpersonal or social function. For instance, systems that suggest, generate, or refine email responses can help users convey positivity (Mieczkowski et al., 2021), politeness, and professionalism in workplace interactions (Miura et al., 2025), while AI-assisted dating profiles can enhance self-presentation (Barkallah et al., 2025). These applications demonstrate that the role of AI extends far beyond grammatical correction to shaping a writer’s tone, style, and communicative intent, thereby aiding in both persuasion and social engagement (Hancock et al., 2020).
This expanding role, however, introduces social friction, particularly when AI’s contribution blurs the boundaries of authorship. Readers often question the authenticity and credibility of AI-generated content. For instance,  Li et al. found that revealing the role of AI in essay writing lowered quality ratings from readers (Li et al., 2024b). Similarly, research on AI-mediated communication showed that online profiles perceived as AI-generated are less trusted (Jakesch et al., 2019). These concerns are particularly salient when writing serves a social purpose, such as persuasion or interaction, where impressions of the author are heavily shaped by perceived authenticity. Recent studies on disclosure dynamics confirm this tension: while AI assistance can improve a text’s fluency, revealing its use can lead readers to judge the author as less sincere or capable (Hwang et al., 2025; Draxler et al., 2024).
AI-assisted writing thus presents a fundamental trade-off. On one hand, it empowers individuals, especially those with language barriers or weaker writing skills, to communicate more effectively (Ningrum and others, 2023; Noy and Zhang, 2023). On the other, the social meaning of authorship complicates its adoption. When readers know or suspect AI involvement, they may perceive the writing—and by extension, the author—as less authentic or trustworthy (Jakesch et al., 2019). Understanding how these perceptions vary across different communicative purposes is therefore critical for designing AI tools that successfully balance expressive power with the need to maintain credibility.
@@ 2.2. Reader Perceptions of AI Disclosure
Research on AI disclosure highlights its paradoxical effect on interpersonal trust. Several studies show that when readers learn AI contributed to a text, their trust in the human author declines even if the quality of the text is unchanged (Jakesch et al., 2019; Lermann Henestrosa and Kimmerle, 2024; Liu et al., 2022; Proksch et al., 2024; Schilke and Reimann, 2025). This “transparency penalty” challenges the conventional assumption that openness builds trust (Sah et al., 2018). The magnitude of this trust erosion is not uniform and appears to be moderated by individual differences. For instance, Schilke and Reimann (2025) found that people with more positive attitudes toward AI show a smaller decline in trust upon disclosure. Furthermore, a user’s AI literacy plays a significant role in shaping their perceptions (Mahmud et al., 2024; Schiavo et al., 2024). AI-literate individuals tend to view its use as pragmatic, whereas less knowledgeable readers may interpret it as a sign of incompetence or laziness (Mahmud et al., 2024; Schiavo et al., 2024).
Complicating this landscape is the fact that people are increasingly unable to distinguish AI-generated text from human writing without explicit disclosure (Májovskỳ et al., 2024; Weber-Wulff et al., 2023), a finding that holds true even for those knowledgeable about large language models (LLMs) (Dugan et al., 2023; Hakam et al., 2024; Porter and Machery, 2024). This creates a risk of misattribution, where readers may credit AI-generated text to a human, only to have their perception negatively altered if AI involvement is later revealed. Despite the importance of this phenomenon, little research has systematically examined how varying degrees of AI involvement influence reader perception across different writing contexts.
Beyond trust, research also reveals mixed attitudes toward the social qualities of AI-generated text. Some studies find that AI generation in communicative scenarios can trigger an “Uncanny Valley effect”, eliciting discomfort (Radivojevic et al., 2024; Yin et al., 2024). Others indicate that AI is perceived as less caring, warm, or empathetic than humans, particularly in sensitive contexts like health communication (Rubin et al., 2025; Wang et al., 2025a). Conversely, other studies show that AI-generated content can be rated as equal or even superior to human-authored text, especially when qualities like efficiency and accuracy are valued (Ayers et al., 2023; Ovsyannikova et al., 2025).
One factor that may explain these divergent findings is people’s prior experience and familiarity with AI. Regular users may view AI authorship as beneficial (Mahmud et al., 2024), while less experienced individuals might perceive it as cold, inauthentic or inadequate (Li et al., 2024b). Together, these findings suggest that AI disclosure has multifaceted effects, influencing not only trust in the author but also perceptions of warmth, competence, and empathy.
@@ 3. Method
To investigate how the disclosure of AI assistance influences reader perception, we conducted a repeated-measures online study. Participants rated their impressions of a text before and after being informed that a specific percentage of the content was generated or edited by AI. Unlike prior work, our primary objective was to examine how perceptions shift across systematically varying levels of disclosed AI contribution. To achieve this, we used a deception-based design: while all texts were entirely AI-generated (with minor human proofreading), participants were told that only specific portions were created or altered by AI. This approach enabled precise control over the perceived degree of AI involvement.
Our study addressed the following research questions:
@@ 3.1. Experiment Design
We grounded our experimental design in the six fundamental acts of writing defined by Berge et al. (Berge et al., 2016): Convince, Interact, Reflect, Describe, Explore, and Imagine. As detailed in Table 1, these acts serve distinct communicative purposes. Convince, Interact, and Reflect are primarily person-oriented, focusing on others, social relationships, and the self, respectively. In contrast, Describe, Explore, and Imagine are object-oriented, centering on the organization, development, and creation of knowledge.
@@
To imagine, narrate, create, and theorize Creation of textual worlds
To create our experimental stimuli, we generated three scenarios for each act, yielding 18 unique texts (Table 2). We created 3–5 initial scenarios per act using GPT-4o [^1] https://chatgpt.com by providing the model with the definition of the act. We then selected the three scenarios we found the most suitable and prompted GPT-4o to generate sentences based on their descriptions. As the study targeted Japanese participants, all texts were presented in Japanese. Pilot testing revealed that sentences generated directly in Japanese by GPT-4o lacked natural fluency. Therefore, we first generated the text in English and then translated it into Japanese. A native Japanese-speaking author reviewed all translated texts, making minor corrections to ensure linguistic quality and naturalness. The final texts averaged 20 sentences and 662 Japanese characters in length. The full English and Japanese scripts are provided as supplementary material.
@@
  A tutorial on operating new equipment
@@
To manipulate the perceived amount of AI involvement, we simulated various levels of AI authorship. For each text, a random proportion of sentences was labeled as generated or edited by AI (in reality, all texts were entirely AI-authored). The distribution of these AI-labeled sentences was randomized using a seed unique to each condition. The disclosed proportion varied from 0% to 100% in 10% increments across conditions, corresponding to approximately two sentences per increment (Figure 1). This design allowed us to model situations where a reader discovers that some portions of a text, from minor local edits to the entire document, were created with AI assistance.
@@ 3.2.1. Pre-disclosure Evaluation
After providing informed consent, participants were each assigned four of the 18 texts, representing different acts of writing, and were asked to read them carefully. For each text, they answered a questionnaire assessing their initial impressions of the (fictitious) author (Figure 1(a)). The questionnaire assessed five perception dimensions using 7-point Likert scales adapted from established instruments (27 statements in total). Trustworthiness (6 statements), Caring/Goodwill (6 statements), and Competence (6 statements) were adapted from McCroskey and Teven (McCroskey and Teven, 1999), and responses ranged from -3 (very negative) to +3 (very positive). Likability (7 statements) and the participant’s Desire for Future Interaction (2 statements), were adapted from Reysen (Reysen, 2005) and Sprecher (Sprecher, 2021). These were rated from -3 (Strongly disagree) to 3 (Strongly agree). All statements are included in Tables 5 and 6 in the Appendix.
The same set of questions was used for both pre- and post-disclosure evaluations, including reverse-coded items and attention checks to ensure data quality.
@@ 3.2.2. Simulated AI Authorship Disclosure and Post-disclosure Evaluation
Following this disclosure, participants answered the same five questions to capture their revised impressions. They were also asked to provide a free-form text response explaining the rationale for any changes in their ratings.
@@ 3.2.3. Post-experimental Questionnaire
After completing evaluations for all four texts, participants filled out a final AI literacy questionnaire to determine whether their familiarity and skills with AI tools might have influenced their perceptions. To measure this competency, we adapted three relevant sub-constructs from the Meta AI Literacy Scale (MAILS) (Carolus et al., 2023) that assess familiarity with and control over AI: “Apply AI”, “Detect AI”, and “AI Persuasion Literacy”. We excluded other MAILS sub-constructs deemed less relevant to the specific scenarios in this study. “Apply AI” measures an individual’s ability to use AI concepts for problem-solving, reflecting proficiency and comfort with AI tools (Ng et al., 2022). “Detect AI” assesses the ability to distinguish AI-enabled technologies from those without AI, a core component of understanding the technology (Long and Magerko, 2020; Wang et al., 2023). Finally, “AI Persuasion Literacy” refers to the ability to recognize and resist AI-driven influence on one’s decisions. Table 6 lists the items for each aspect of AI literacy.
@@
Figure 1. The user interfaces of the experiment: (a) pre-rating and (b) post-rating phases. The original interface was in Japanese and the text shown here is translated for presentation, but the layout and visual design are preserved.
A two-panel figure. Part (a) shows the pre-disclosure interface with a block of text on the left and a questionnaire on the right. Part (b) shows the post-disclosure interface, which is identical except that several sentences in the left text panel are highlighted in red to indicate AI generation.